OpenStudy (anonymous):

I have a couple of 8th grade math questions. Help?

OpenStudy (anonymous):

Sure i can try

OpenStudy (anonymous):

Which description best describes the solution to the following system of equations? y = −one halfx + 9 y = x + 7 Lines y = −one halfx + 9 and y = x + 7 intersect the x-axis. Lines y = −one halfx + 9 and y = x + 7 intersect the y-axis. Line y = −one halfx + 9 intersects the origin. Line y = −one halfx + 9 intersects line y = x + 7.

OpenStudy (anonymous):

Solution to the system of equations means the common points which satisfy both the equations simultaneously

OpenStudy (anonymous):